A rock is thrown straight down with initial velocity of 14, 5 m/s from a cliff .what is the rocks displacement after 2, 0 s


1
Expert's answer
2022-04-17T17:06:34-0400
"d=v_0t+at^2/2\\
=14.5*2.0+9.8*2.0^2/2=48.6\:\rm m"